OpenStudy (anonymous):

Simplify the radical expression.. square root of 128

OpenStudy (anonymous):

Well is there a square root within 128...

OpenStudy (anonymous):

\[\sqrt{2*64}\]\[\sqrt{2}\sqrt{64}\]\[8\sqrt{2}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You are surely welcome!

OpenStudy (anonymous):

what about square root of 500?

OpenStudy (anonymous):

Use your calculator to see if there is a square embedded in 500... that is, what are the 2 [or more] highest factors of 5 that multiply to give you 500...

OpenStudy (anonymous):

So you can go 500/1 500/2, 500/3... which would take forever or you can use the factors 25, 5, 4, etc...

OpenStudy (anonymous):

Let me know if you get it..

OpenStudy (anonymous):

so 5 * 100?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So what's your final answer?

OpenStudy (anonymous):

5 and 20? or 10?

OpenStudy (anonymous):

\[10\sqrt{5}\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Do you understand how I got this?

OpenStudy (anonymous):

\[\sqrt{100}\sqrt{5}\]

OpenStudy (anonymous):

bc 10 will go into 100 10 times?

OpenStudy (anonymous):

no because the square root of 100 is 10

OpenStudy (anonymous):

Are you using your scientific calculator to do these questions?

OpenStudy (anonymous):

or do you have to do them out of your head?

OpenStudy (anonymous):

No i dont have one

OpenStudy (anonymous):

google online scientific calculator

OpenStudy (anonymous):

ok thanks a bunch!